OpenStudy (anonymous):

A person who is 1.5 m tall casts a shadow of 8 m. The distance along the ground from the person (N) to the flagpole (G) is 32 m. Find the height of the flagpole (FG), showing the calculation that led to your answer.

OpenStudy (anonymous):

Hi Welcome to Openstudy @LillianFaith011 :) First, let's visualize the problem by drawing a diagram, then apply similar triangles. 

OpenStudy (anonymous):

alright :) it will be the same thing \(\sf \Large \frac{person's\ height}{shadow}=\frac{flagpole's\ height}{distance}\) \(\sf \Large \frac{1.5}{8}=\frac{flagpole's\ height}{32}\) cross multiply and solve for the flagpole's height

OpenStudy (anonymous):

Thank you so Much!!!!!!!!!

OpenStudy (anonymous):

did you get the answer already? ^_^

OpenStudy (anonymous):

and no problem, if you have more questions, just make another post and other users can also help you :) goodluck with studies

OpenStudy (anonymous):

Yes, I couldnt figure out what to do.. But once i got to cross multiply it was easy! THank you

OpenStudy (anonymous):

:)